A college student invested part of a $25,000 inheritance at 7% interest and the rest at 6%. If his annual interest is $1,671 how much did he invest at 6%? If I told you the answer is $8,000, in your own words, using complete sentences, explain how you would solve the problem.

Answer by checkley77(12844)

I WOULD USE THE STANDARD INVESTMENT FORMULA FOR A 2 PART INVESTMENT
USING 2 DIFFERENT INTEREST FIGURES.
WITH THE $8,000 @ 6% PLUS THE REMAINING $17,000 @ 7% EQUAL TO THE TOTAL INTERST OF $1,671.
THUS:
.07(25,000-8,000)+.06*8,000=1,671
.07*17,000+480=1,671
1,190+480=1,671
1,670~1,671